# Identifying Policy Gradient Subspaces

### v1

@@ 1 Introduction
Deep reinforcement learning (RL) has marked significant achievements in numerous challenging problems, ranging from Atari games (Mnih et al., 2013) to various real robotic challenges, such as contact-rich manipulation (Gu et al., 2017; Kalashnikov et al., 2018), complex planning problems (Everett et al., 2018; Ao et al., 2022), and hard-to-control dynamic tasks (Cheng et al., 2023; Kaufmann et al., 2023). Despite these notable successes, deep RL methods are often brittle due to the use of function approximators with a multitude of parameters and persistently changing data distributions – a setting notoriously hard for optimization. Deep RL, in its vanilla form, operates under the premise of limited prior knowledge and structural information about the problem and instead resorts to learning from experience by directly interacting with its environment.
@@
Recent advances in supervised learning (SL) introduce more structured optimization procedures by identifying and harnessing gradient subspaces (Gur-Ari et al., 2018). These approaches reveal that the gradients utilized for neural network optimization reside in a low-dimensional, slowly-changing subspace (Vinyals & Povey, 2012; Li et al., 2022a; Gauch et al., 2022; Duda, 2019). Exploiting this structure has garnered attention since it enables the optimization to be carried out in this reduced-dimensional subspace, yielding enhanced efficiency with minimal, if any, loss in performance.
@@
Despite the promise of subspaces in SL, their adoption in deep RL has remained limited. A straightforward way to transfer these principles is to find lower-dimensional subspaces in policy gradient approaches (Peters & Schaal, 2008). Policy gradient (PG) methods directly estimate the gradient of the RL objective to update the policy’s parameters using some form of stochastic gradient descent (SGD). Since most SL approaches using subspaces enter at the level of the SGD optimization, PG algorithms would be a natural choice to leverage the knowledge about subspaces from SL in the RL context. Nevertheless, in RL, subspaces have primarily been explored within the realm of evolutionary strategies (Maheswaranathan et al., 2019; Li et al., 2018), representation learning (Le Lan et al., 2023), and transfer learning (Gaya et al., 2022). A possible explanation is the constantly changing data distribution of RL due to continual exploration that seems to hinder the identification of gradient subspaces. The limited body of studies using subspaces in PG algorithms underlines the need for a more profound discussion in this domain.
@@
This paper conducts a comprehensive empirical evaluation of gradient subspaces in the context of PG algorithms, assessing their properties across various simulated RL benchmarks. Our experiments reveal several key findings: (i) there exist parameter-space directions that exhibit significantly larger curvature compared to the other parameter-space directions, (ii) the gradients live in the subspace spanned by these directions, and (iii) the subspace remains relatively stable throughout the RL training. Additionally, we analyze the gradients of the critic – an integral part of the PG estimation in actor-critic methods – and observe that the critic subspace often exhibits less variability and retains a larger portion of its gradient compared to the actor subspace. We also test the robustness of PG subspaces regarding mini-batch approximations of the gradient that are used in practice during training and evaluate a similar mini-batch approximation of the Hessian. Lastly, we explore the extent to which the variation in the data distribution influences the aforementioned subspace analysis by conducting the experiments with both an on-policy as well as an off-policy algorithm, the latter of which reuses previously collected data for training.
@@
By shedding light on subspaces in deep RL, this paper aims to provide valuable insights that can potentially enhance RL performance by advancing parameter-space exploration or enabling second-order optimization. We begin by reviewing existing literature on subspace approaches in Section 2, followed by a recapitulation of the RL preliminaries in Section 3 to better understand the analysis of gradient subspaces in RL in Section 4. Section 5 concludes this work with a discussion of the results and implications of our work. The code for our experiments is available on the project website.
@@ 2 Related work
Numerous works have studied the use of gradient subspaces in SL. These works can be roughly divided into informed and random subspace approaches.
Informed gradient subspaces in supervised learning
@@
Informed methods identify the subspace from some part of the training process. For instance, Gur-Ari et al. (2018) show that the gradients used for neural network optimization lie in a low-dimensional subspace of the parameter-space directions with the highest curvature in the loss. Furthermore, they show that this subspace changes slowly throughout the training. Recent work has highlighted numerous benefits of these insights. Li et al. (2022a) apply principal component analysis (PCA) on the network parameters at previous checkpoints and use the top principal components as subspace. They apply SGD and BFGS algorithms in the subspace and demonstrate superior learning performance compared to training in the original parameter space. Similarly, Tuddenham et al. (2020) propose to use a BFGS method in a low-dimensional subspace defined by the full-batch gradients with respect to the individual classification labels. They apply their method to classification and use the set of full-batch gradients for samples from each individual class as gradient subspace. Sohl-Dickstein et al. (2014) construct a second-order Taylor approximation of individual loss terms efficiently in a low-dimensional subspace that captures recently observed gradients. With this approximation, the overall loss is a sum of quadratic functions that can be optimized in closed form. Chen et al. (2022) utilize insights about the low-dimensional subspaces to learn an optimizer that scales to high-dimensional parameter spaces. Traditionally, learned optimizers are limited in scale by the high output dimensionality of the network. By outputting parameter updates in the low-dimensional subspace, Chen et al. circumvent this challenge and efficiently learn to optimize large neural networks. Gauch et al. (2022) apply the knowledge of low-dimensional gradient subspace to few-shot adaptation by identifying a suitable subspace on the training task and utilizing it for efficient finetuning on the test task. Likewise, Zhou et al. (2020) tackle differential privacy by identifying a gradient subspace on a public dataset and projecting gradients calculated from the private dataset into this subspace. Adding noise in the subspace instead of the parameter space lowers the error bound to be linear in the number of subspace dimensions instead of the parameter dimensions. Li et al. (2022b) tackle an overfitting problem of adversarial training by identifying a gradient subspace in an early training phase before overfitting occurs. By projecting later gradients into this subspace, they enforce training dynamics similar to the early training phase, which mitigates the overfitting problem.
Random gradient subspaces in supervised learning
@@
Gressmann et al. (2020) describe different modifications to improve neural network training in random subspaces, like resampling the subspace regularly and splitting the network into multiple components with individual subspaces. They utilize these subspaces to reduce the communication overhead in a distributed learning setting by sending the parameter updates as low-dimensional subspace coefficients. Li et al. (2018) quantify the difficulty of a learning task by the dimensionality of the subspace needed to reach good performance. Larsen et al. (2021) compare informed and random subspaces and find that with informed subspaces, typically significantly fewer dimensions are needed for training. Furthermore, they note the benefit of first training in the original space for some time before starting subspace training, as this increases the probability of the subspace intersecting low-loss regions in the parameter space.
@@
Only a handful of works use subspaces for RL. Li et al. (2018) sample random subspaces of the parameter space to find the lowest dimensional subspace without significant loss in performance. To that end, the authors optimize the value function of deep Q-networks (DQN) (Mnih et al., 2013) as well as the population distribution -– a distribution over the policy parameters –- of evolutionary strategies (ES) (Salimans et al., 2017), using the subspace-adapted version of SGD. Maheswaranathan et al. (2019) define a search distribution for ES that is elongated along an informed subspace spanned by surrogate gradients. Despite not being an RL approach (but optimizing over policies within the ES framework), this work still reveals the important insight that the policy is only indirectly affected in these scenarios. Gaya et al. (2022) enable improved generalization in online RL by learning a subspace of policies. Recently, they extended their method to scale favorably to continual learning settings (Gaya et al., 2023). Both works consider the subspace of policy parameters rather than the subspace of policy gradients.
@@ 3.2.1 Proximal Policy Optimization
On-policy PG methods typically optimize the policy via an objective such as with and being an estimator of the advantage function at timestep and denotes the policy before the update (Kakade & Langford, 2002). However, optimizing this objective can result in excessively large updates, leading to instabilities and possibly divergence. Proximal Policy Optimization (PPO) (Schulman et al., 2017) is an on-policy actor-critic algorithm designed to address this issue by clipping the probability ratio to the interval , which removes the incentive for moving outside the interval, resulting in the following actor loss.
@@ 4 Gradient subspaces in policy gradient algorithms
In Section 2, we have highlighted several works from SL that utilize low-dimensional, slowly-changing gradient subspaces for improving the learning performance or as an analysis tool. Naturally, we would like to transfer these benefits to policy gradient algorithms. The training in RL is significantly less stationary than in the supervised setting (Bjorck et al., 2022). As the RL agent changes, the data distribution shifts since the agent continuously interacts with its environment when collecting the data. Furthermore, the value of a state also depends on the agent’s behavior in future states. Thus, the targets for the actor and critic networks change constantly. These crucial differences between SL and RL underscore the need to analyze to which extent insights about gradient subspaces transfer between these settings.
@@
The analysis presented in this section focuses on two policy gradient algorithms: PPO (Schulman et al., 2017) and SAC (Haarnoja et al., 2018), which are popular instantiations of on-policy and off-policy RL. We apply the algorithms to twelve benchmark tasks from OpenAI Gym (Brockman et al., 2016), Gym Robotics (Plappert et al., 2018a), and the DeepMind Control (DMC) Suite (Tunyasuvunakool et al., 2020). Our code builds upon the algorithm implementations of Stable Baselines3 (Raffin et al., 2021). The learning curves are displayed in Appendix B. We ran each experiment for 10 random seeds and plot the mean and standard deviation for all results Sections 4.2 and 4.3. Due to space constraints, in the following sections, we only show analysis results for selected tasks. Detailed results for all twelve tasks can be found in Appendix C. Moreover, we conduct an evaluation of the impact of the RL algorithm’s hyperparameters on the gradient subspace in Appendix D.
@@
For the following analyses, we calculate Hessian eigenvectors of the loss with respect to the network parameters via the Lanczos method (Lehoucq et al., 1998) since it is an efficient method for estimating the top eigenvectors that avoids explicitly constructing the Hessian matrix. Since we can only estimate the Hessian from data, we use a large number of state-action pairs to obtain precise estimates of the eigenvectors of the true Hessian, similar to how Ilyas et al. (2020) estimate the true policy gradient. For PPO, we collect on-policy samples. This would, however, not be faithful to the diverse distribution of off-policy data that SAC uses for training. To match this data distribution for the analysis, we save the replay buffer during training and use the data of the complete replay buffer for estimating the Hessian. Note that the replay buffer also has a capacity of samples but is not completely filled at the beginning of training.
@@
As mentioned in Section 3, SAC and PPO each train two different networks, an actor and a critic. We, therefore, conduct our analysis for each network individually. To verify that there exist directions with high curvature spanning a subspace that stays relatively stable throughout the training and that contains the gradient, we check three conditions:
@@
There exist parameter-space directions of significantly larger curvature in the actor/critic loss than the other directions.
@@ 4.1 The loss curvature is large in a few parameter-space directions
The Hessian matrix describes the curvature of a function, with the eigenvectors being the directions of maximum and minimum curvature. The corresponding eigenvalues describe the magnitude of the curvature along these directions. Therefore, we verify condition i) by plotting the spectrum of Hessian eigenvalues for the actor and critic loss of PPO relative to the respective model parameters in Figure 1. The plots show that there are a few large eigenvalues for both the actor and critic loss. All remaining eigenvalues are distributed close to zero. These plots confirm that there are a few directions with significantly larger curvature; in other words, the problem is ill-conditioned.
@@ 4.2 The gradient lies in the high-curvature subspace
To verify condition ii) that the high-curvature subspace contains the gradients of the respective loss, we measure how much of these gradients can be represented in the subspace. Let be the semi-orthogonal matrix that projects into the high-curvature subspace. consists row-wise of the largest Hessian eigenvectors. We compute the relative projection error, i.e., the relative difference between the original gradient and the projected gradient that is the result of mapping into the high-curvature subspace and back into the original space. The fraction of the gradient that can be represented in the subspace is given by
@@
We derive this equality in Appendix A. Note that holds, where implies that the subspace perfectly contains the gradient, while means that the gradient lies entirely outside the subspace. Due to the normalization by , this criterion is invariant to the scale of the gradient, which enables comparing gradient subspaces of different models and models at different stages of the training.
To evaluate how the gradient subspace fraction evolves over time, we evaluate the gradient subspace fraction at checkpoints every steps during the RL training. To compactly visualize this data, we split the training into three phases: initial, training, and convergence, and for each phase, we average the results of all timesteps of that phase. Since the algorithms require a different number of timesteps for solving each of the environments and reach different performance levels, we define the following heuristic criterion for the training phases. We first smooth the learning curves by averaging over a sliding window and compute the maximum episode return over the smoothed curve. Next, we calculate the improvement relative to the episode return of the initial policy at each timestep of the smoothed learning curve as
We define the end of the initial phase as the timestep when the agent reaches 10% of the total improvement, i.e., , for the first time. Similarly, we define the start of the convergence phase as the first timestep at which it reaches 90% of the total improvement, i.e., .
We choose as subspace dimensionality since this subspace already largely captures the gradients, and the first eigenvectors can still be calculated reasonably efficiently with the Lanczos method. Appendix C displays results for different subspace sizes. With the tuned hyperparameters from RL Baselines3 Zoo that we use for training, the PPO actor and critic usually contain around parameters, and the SAC actor and critic around and parameters (2 Q-networks à parameters), respectively. Hence, the subspace dimensionality is around 2% the size of the parameters for PPO and around 0.14% and 0.07% for SAC.
@@
Figure 2 shows the value of the gradient subspace fraction for PPO and SAC on four different tasks, divided into the three training phases. Note that for an uninformed random projection, the gradient subspace fraction would be in expectation, i.e., the ratio of the original and subspace dimensionalities ( for PPO and / for SAC). The results in Figure 2 show a significantly higher gradient subspace fraction, which means that the gradients computed by PPO and SAC lie to a large extent in the high-curvature subspace. For PPO, we observe that the fraction of the gradient in the subspace is considerably higher for the critic than for the actor. Furthermore, the gradient subspace fraction is also often markedly higher for SAC than for PPO. This finding is particularly significant since the subspace size corresponds to a significantly lower percentage of the parameter dimensionality for SAC than for PPO. We hypothesize that the effect is caused by the off-policy nature of SAC. In the off-policy setting, the training distribution for the networks changes slowly since the optimization reuses previous data. In this regard, SAC is closer than PPO to the supervised learning setting, where the data distribution is fixed and for which Gur-Ari et al. (2018) report high gradient subspace fractions. Still, the subspace fraction for PPO is significant, considering that the dimensionality of the subspace is merely 2% of the original parameter space. Furthermore, for PPO, the subspace fraction often improves after the initial phase. Similarly, Gur-Ari et al. (2018) report for the supervised learning setting that the gradient moves into the subspace only after some initial steps.
Moreover, the precise gradient, computed with a large number of samples, tends to lie slightly better in the subspace than the mini-batch gradient. The noise resulting from the low-sample approximation seems to slightly push the gradient out of the subspace. However, since the difference is typically small, the gradient subspace is still valid for the low-sample gradient estimates used during RL training. Lastly, even the subspace identified with the mini-batch Hessian captures the gradient to a significant extent. This property is crucial since it implies that we do not need access to the precise Hessian, which is costly to compute and might require additional data. Instead, we can already obtain a reasonable gradient subspace from the mini-batch Hessian.

### v3

@@ 1 Introduction
Deep reinforcement learning (RL) has marked significant achievements in numerous challenging problems, ranging from Atari games (Mnih et al., 2013) to various real robotic challenges, such as contact-rich manipulation (Gu et al., 2017; Kalashnikov et al., 2018), complex planning problems (Everett et al., 2018; Ao et al., 2022), and hard-to-control dynamic tasks (Cheng et al., 2023; Kaufmann et al., 2023). Despite these notable successes, deep RL methods are often brittle due to the use of function approximators with large numbers of parameters and persistently changing data distributions – a setting notoriously hard for optimization. Deep RL, in its vanilla form, operates under limited prior knowledge and structural information about the problem, consequently requiring large numbers of interactions with the environment to reach good performance.
@@
For supervised learning (SL), Gur-Ari et al. (2018) demonstrated that the gradients utilized for neural network optimization reside in a low-dimensional, slowly-changing subspace. Based on this insight, recent works introduce more structured optimization procedures for SL by identifying and harnessing these gradient subspaces. Exploiting this structure enables the optimization to be carried out in a reduced-dimensional subspace, yielding enhanced efficiency with minimal, if any, loss in performance (Li et al., 2018; Gressmann et al., 2020; Larsen et al., 2021; Li et al., 2022a).
@@
Despite the benefits of subspace methods in SL, their adoption in deep RL has remained limited. A straightforward way to transfer these principles is to find lower-dimensional subspaces in policy gradient approaches (Peters & Schaal, 2008). Policy gradient (PG) methods estimate the gradient of the RL objective to update the policy’s parameters using some form of stochastic gradient descent (SGD). Since most SL approaches using subspaces operate at the level of the SGD optimization, PG algorithms would be a natural choice to leverage the knowledge about subspaces from SL in the RL context. Nevertheless, in RL, such methods have been explored primarily within the realm of evolutionary strategies (Maheswaranathan et al., 2019), representation learning (Le Lan et al., 2023), and transfer learning (Gaya et al., 2022). A possible explanation is the constantly changing data distribution of RL due to continual exploration that intuitively seems to hinder the identification of gradient subspaces. The limited body of studies using subspaces in PG algorithms underlines the need for a more profound discussion in this domain.
@@
This paper conducts a comprehensive empirical evaluation of gradient subspaces in the context of PG algorithms, assessing their properties across various simulated RL benchmarks. Our experiments reveal several key findings: (i) there exist parameter-space directions that exhibit significantly larger curvature compared to other parameter-space directions, (ii) the gradients live in the subspace spanned by these directions, and (iii) the subspace remains relatively stable throughout the RL training. Additionally, we analyze the gradients of the critic – an integral part of the PG estimation in actor-critic methods – and observe that the critic subspace often exhibits less variability and retains a larger portion of its gradient compared to the actor subspace. We also test the robustness of PG subspaces regarding mini-batch approximations of the gradient that are used in practice during training and evaluate a similar mini-batch approximation of the Hessian. Lastly, we explore the extent to which the variation in the data distribution influences the aforementioned subspace analysis by conducting experiments with both an on-policy as well as an off-policy algorithm, the latter of which reuses previously collected data for training.
@@
By shedding light on gradient subspaces in deep RL, this paper provides insights that can potentially enhance RL performance by advancing parameter-space exploration or enabling second-order optimization. We begin by reviewing existing literature on subspace approaches in Section 2, followed by a recapitulation of the RL preliminaries in Section 3 as a foundation for the analysis of gradient subspaces in RL in Section 4. Section 5 concludes this work with a discussion of the results and implications of our work. The code for our experiments is available on the project website.
@@ 2 Related work
Numerous works have studied the use of gradient subspaces in SL. These works can be roughly divided into informed and random subspace approaches. In the following, we give an overview of these papers and highlight works that investigate related concepts in RL.
Informed subspaces in supervised learning
@@
Informed subspace methods identify the subspace from some part of the training process. For instance, Gur-Ari et al. (2018) show that the gradients used for neural network optimization lie in a low-dimensional subspace of the parameter-space directions with the highest curvature in the loss. Furthermore, they show that this subspace changes slowly throughout the training. Recent works have highlighted numerous benefits of these insights. Li et al. (2022a) apply principal component analysis on the network parameters at previous checkpoints and use the top principal components as subspace. They apply SGD and BFGS algorithms in the subspace and demonstrate superior learning performance compared to training in the original parameter space. Similarly, Tuddenham et al. (2020) propose to use a BFGS method in a low-dimensional subspace defined by the full-batch gradients with respect to individual classification labels. Sohl-Dickstein et al. (2014) construct a second-order Taylor approximation of individual loss terms efficiently in a low-dimensional subspace that captures recently observed gradients. With this approximation, the overall loss is a sum of quadratic functions that can be optimized in closed form. Chen et al. (2022) utilize insights about the low-dimensional subspaces to learn an optimizer that scales to high-dimensional parameter spaces. Traditionally, learned optimizers are limited in scale by the high output dimensionality of the network. By outputting parameter updates in the low-dimensional subspace, Chen et al. circumvent this challenge and efficiently learn to optimize large neural networks. Gauch et al. (2022) apply the knowledge of low-dimensional gradient subspace to few-shot adaptation by identifying a suitable subspace on the training task and utilizing it for efficient finetuning on the test task. Likewise, Zhou et al. (2020) tackle differential privacy by identifying a gradient subspace on a public dataset and projecting gradients calculated from the private dataset into this subspace. Adding noise in the subspace instead of the parameter space lowers the error bound of differential privacy to be linear in the number of subspace dimensions instead of the parameter dimensions. Li et al. (2022b) tackle an overfitting problem of adversarial training by identifying a gradient subspace in an early training phase before overfitting occurs. By projecting later gradients into this subspace, they enforce that the training dynamics remain similar to the early training phase, which mitigates the overfitting problem.
Random subspaces in supervised learning
@@
Gressmann et al. (2020) describe different modifications to improve neural network training in random subspaces, like resampling the subspace regularly and splitting the network into multiple components with individual subspaces. They utilize these subspaces to reduce the communication overhead in a distributed learning setting by sending the parameter updates as low-dimensional subspace coefficients. Li et al. (2018) quantify the difficulty of a learning task by its intrinsic dimensionality, i.e., the dimensionality of the random subspace needed to reach good performance. Aghajanyan et al. (2021) apply a similar analysis to language models and find that the effectiveness of fine-tuning language models is tied to a low intrinsic dimensionality. Based on these findings, Hu et al. (2021) propose a method for fine-tuning large models efficiently via low-rank parameter updates. Larsen et al. (2021) compare informed and random subspaces and find that with informed subspaces, typically significantly fewer dimensions are needed for training. Furthermore, they note the benefit of first training in the original space for some time before starting subspace training, as this increases the probability of the subspace intersecting with low-loss regions in the parameter space.
@@
Only a handful of works use subspace methods for RL. Tang et al. (2023) identify a parameter subspace through singular value decomposition of the policy parameters’ temporal evolution. They propose to cut off the components of the update step that lie outside this subspace and additionally elongate the update in the directions corresponding to the largest singular values. In contrast to our work, they focus solely on the policy network, consider only off-policy learning algorithms, and do not establish a connection to the curvature of the learning objective. Li et al. (2018) quantify the difficulty of RL tasks by their intrinsic dimensionality. To that end, the authors optimize the value function of Deep Q-Networks (DQN) (Mnih et al., 2013) as well as the population distribution – a distribution over the policy parameters – of evolutionary strategies (ES) (Salimans et al., 2017) in random subspaces. Maheswaranathan et al. (2019) define a search distribution for ES that is elongated along an informed subspace spanned by surrogate gradients. While the authors do not evaluate their approach for RL, they highlight the relevance to RL due to the method’s ability to use surrogate gradients generated from learned models, similar to the use of the critic in actor-critic methods. Gaya et al. (2022) enable improved generalization in online RL by learning a subspace of policies. Recently, they extended their method to scale favorably to continual learning settings (Gaya et al., 2023). Both methods explicitly learn a subspace of policy parameters, while our work investigates the natural training dynamics of PG methods.
@@ 3.2.1 Proximal Policy Optimization
On-policy PG methods typically optimize the policy via an objective such as with and being an estimator of the advantage function at timestep and denoting the policy before the update (Kakade & Langford, 2002). However, optimizing this objective can result in excessively large updates, leading to instabilities and possibly divergence. Proximal Policy Optimization (PPO) (Schulman et al., 2017) is an on-policy actor-critic algorithm designed to address this issue by clipping the probability ratio to the interval , which removes the incentive for moving outside the interval, resulting in the following actor loss.
@@ 4 Gradient subspaces in policy gradient algorithms
In Section 2, we have highlighted several works from SL that utilize low-dimensional gradient subspaces for improving the learning performance. Naturally, we would like to transfer these benefits to policy gradient algorithms. However, the training in RL is significantly less stationary than in the supervised setting (Bjorck et al., 2022). As the RL agent changes, the data distribution shifts since the data is generated by the agent’s interactions with its environment. Furthermore, the value of a state also depends on the agent’s behavior in future states. Thus, the targets for the actor and critic networks change constantly. These crucial differences between SL and RL underscore the need to analyze to which extent insights about gradient subspaces transfer between these settings.
@@
The analysis presented in this section focuses on two policy gradient algorithms: PPO (Schulman et al., 2017) and SAC (Haarnoja et al., 2018), which are popular instantiations of on-policy and off-policy RL. We apply the algorithms to twelve benchmark tasks from OpenAI Gym (Brockman et al., 2016), Gym Robotics (Plappert et al., 2018a), and the DeepMind Control Suite (Tunyasuvunakool et al., 2020). Our code builds upon the algorithm implementations of Stable Baselines3 (Raffin et al., 2021). The learning curves are displayed in Appendix A. We ran each experiment for 10 random seeds and plot the mean and standard deviation for the results in Sections 4.2 and 4.3. Due to space constraints, we show the analysis results only for selected tasks and present detailed results for all twelve tasks in Appendix B. Moreover, we conduct an evaluation of the impact of the RL algorithm’s hyperparameters on the gradient subspace in LABEL:app:impact_of_suboptimal_hyperparameters.
@@
For the following analyses, we calculate Hessian eigenvectors of the loss with respect to the network parameters via the Lanczos method (Lehoucq et al., 1998) since it is an efficient method for estimating the top eigenvectors that avoids explicitly constructing the Hessian matrix. Since we can only estimate the Hessian from data, we use a large number of state-action pairs to obtain precise estimates for the eigenvectors of the true Hessian, similar to how Ilyas et al. (2020) approximate the true policy gradient. For PPO, we collect on-policy samples. This would, however, not be faithful to the diverse distribution of off-policy data that SAC uses for training. To match this data distribution for the analysis, we save the replay buffer during training and use the data of the complete replay buffer for estimating the Hessian. Note that the replay buffer also has a capacity of samples but is not completely filled at the beginning of training.
@@
As mentioned in Section 3, SAC and PPO each train two different networks, an actor and a critic. We, therefore, conduct our analysis for each network individually. To verify that there exist high-curvature directions spanning a subspace that stays relatively stable throughout the training and that contains the gradient, we check three conditions:
@@
Some parameter-space directions exhibit significantly larger curvature in the actor/critic loss than other directions.
@@ 4.1 The loss curvature is large in a few parameter-space directions
The Hessian matrix describes the curvature of a function, with the eigenvectors being the directions of maximum and minimum curvature. The corresponding eigenvalues describe the magnitude of the curvature along these directions. Therefore, we verify condition i) by plotting the spectrum of Hessian eigenvalues for the actor and critic loss of PPO with respect to the network parameters in Figure 1. The plots show that there are a few large eigenvalues for both the actor and critic loss. All remaining eigenvalues are distributed close to zero. These plots confirm that there are a few directions with significantly larger curvature; in other words, the problem is ill-conditioned.
@@ 4.2 The gradient lies in the high-curvature subspace
To verify condition ii) that the high-curvature subspace contains the gradients of the respective loss, we measure how well these gradients can be represented in the subspace. Let be the semi-orthogonal matrix that projects into the high-curvature subspace. consists row-wise of the largest Hessian eigenvectors. We compute the relative projection error, i.e., the relative difference between the original gradient and the projected gradient that is the result of mapping into the high-curvature subspace and back into the original space. The fraction of the gradient that can be represented in the subspace is then given by
@@
We derive this equality in LABEL:app:derivation_gradient_subspace_fraction. Note that holds, where implies that the subspace perfectly contains the gradient, while means that the gradient lies entirely outside the subspace. Due to the normalization by , this criterion is invariant to the scale of the gradient, which enables comparing gradient subspaces of different models and models at different stages of the training.
To evaluate how the gradient subspace fraction evolves over time, we evaluate the criterion at checkpoints every steps during the RL training. To compactly visualize this data, we split the training into three phases: initial, training, and convergence, and for each phase, we average the results of all timesteps of that phase. Since the algorithms require different numbers of timesteps for solving each of the tasks and reach different performance levels, we define the following heuristic criterion for the training phases. We first smooth the learning curves by averaging over a sliding window and compute the maximum episode return over the smoothed curve. Next, we calculate the improvement relative to the episode return of the initial policy at each timestep of the smoothed learning curve as
We then define the end of the initial phase as the first timestep at which the agent reaches 10% of the total improvement, i.e., . Similarly, we define the start of the convergence phase as the first timestep at which the agent reaches 90% of the total improvement, i.e., .
We choose as subspace dimensionality since this subspace already largely captures the gradients, and the largest eigenvectors can still be calculated reasonably efficiently with the Lanczos method. Appendix B displays results for different subspace sizes. With the tuned hyperparameters from RL Baselines3 Zoo that we use for training, the PPO actor and critic usually contain around parameters, and the SAC actor and critic around and parameters (2 Q-networks à parameters), respectively. Hence, the subspace dimensionality is around 2% the size of the parameters for PPO and around 0.14% and 0.07% for SAC.
@@
Figure 2 shows the value of the gradient subspace fraction for PPO and SAC on four different tasks, divided into the three training phases. Note that for an uninformed random projection, the gradient subspace fraction would be in expectation, i.e., the ratio of the original and subspace dimensionalities ( for PPO and for SAC’s actor and for its critic). The results in Figure 2 show a significantly higher gradient subspace fraction, which means that the gradients computed by PPO and SAC lie to a large extent in the high-curvature subspace. We observe that the fraction of the gradient in the subspace is considerably higher for the critic than for the actor. Furthermore, the gradient subspace fraction is also often higher for SAC’s actor than for PPO’s. This finding is particularly significant since the subspace size corresponds to a significantly lower percentage of the parameter dimensionality for SAC than for PPO. We hypothesize that the effect is caused by the off-policy nature of SAC. In the off-policy setting, the training distribution for the networks changes slowly since the optimization reuses previous data. In this regard, SAC is closer than PPO to the supervised learning setting, where the data distribution is fixed and for which Gur-Ari et al. (2018) report high gradient subspace fractions. Still, the subspace fraction for PPO is significant, considering that the dimensionality of the subspace is merely 2% of the original parameter space. Furthermore, for PPO, the subspace fraction often improves after the initial phase. Similarly, Gur-Ari et al. (2018) report for the supervised learning setting that the gradient starts evolving in the subspace only after some initial steps. However, for the SAC actor, this trend appears to be reversed, with the gradient subspace fraction being highest in the initial steps.
Moreover, the precise gradient, computed with a large number of samples, tends to lie better in the subspace than the mini-batch gradient. The noise resulting from the low-sample approximation seems to perturb the gradient out of the subspace. However, since the difference is typically small, the gradient subspace is still valid for the low-sample gradient estimates used during RL training. Lastly, even the subspace identified with the mini-batch Hessian captures the gradient to a significant extent. This property is crucial since it implies that we do not need access to the precise Hessian, which is costly to compute and might require additional data. Instead, we can already obtain a reasonable gradient subspace from the mini-batch Hessian.